OpenStudy (anonymous):

2x+3=-4x-9y 2x+8y=3-4x can i solve this question with substitution method

OpenStudy (anonymous):

2x+4x=-9y-3 2x+4x=-8y+3 - - = + - 0 = -y-6 6=-y y=-6 m i going right

OpenStudy (anonymous):

That's not the substitution method; that's subtraction. Because you solved them both to equal 2x+4x, you could substitute them by taking the right-hand sides of both of the equations and setting them equal to each other:\[-9y-3=-8y+3\]Then solve for y:\[9y=8y-6\]\[y=-6\]

OpenStudy (anonymous):